OpenStudy (anonymous):

The National Association for Women in Science asked recent high school grads if they had taken certain science classes. Of those surveyed, 68 said they had taken physics, 108 said they had taken chemistry and 12 said they had taken both. Nine said they had taken neither. How many recent high school grads were surveyed?

OpenStudy (anonymous):

12 people said yes twice. That's 24, 12 too many.

OpenStudy (anonymous):

197, right? all you need to do is add all the numbers. it never said the ones in the both categories were added to the individual physics and chemistry numbers.

OpenStudy (anonymous):

Thats what i thought at first but my book is showing that you have to subtract the twelve from both physics and chemistry, so I am really not too sure.. was just wondering if anybody else comes up with what i got which is 173

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

so it is 173?

OpenStudy (anonymous):

guess your gone now.. thanks for yourhelp..

OpenStudy (anonymous):

173

OpenStudy (anonymous):

108 had taken chemistry but 12 of them also has physics so the group of only chemistry is 108-12=96 68had taken physics but 12 of them also has chemistry so the group of only physics is 68-12=56 And you the 12 people with both so 96+56+12=164

OpenStudy (anonymous):

+9 = 173